19. In 1995, there were 33.8 million cellular subscribers in the U.S.In 1999 there were 86 million.
a. Let x = 5 represent 1995 and x = 9 represent 1999. Find the average rate of change in the number of subscribers in the US.
b. Write a linear function that models the number of cellular subscribers in the US.
c. Interpret the average rate of change and the y-intercept using a complete sentence and be specific to this problem.
d. According to your calculations, how many cellular subscribers will there be in2012?
e. Is this a good estimation? Why or why not.

Answer by ewatrrr(24785)

(9, 86 )
(5, 33.8) rate of change = m = 52.2/4 = 13.05
.........
y -86 = 13.05(x - 9)
y = 13.05x -31.45
..........
rate of growth is ~13M /year since the 1990 total of 31.45M
........
x = 12
y = 13.05(12) - 31.45 = 125.15 million
assuming rate of growth remains stable, Yes.